6. Identify the significance of standard deviation in data analytics using Table 5A data as a reference.

Data Table 5A

Standard Deviation Table

Coder Status

Mean

N

Standard Deviation

Inpatient Coding Specialist II

93.0000

3

5.0000

Inpatient Coding Specialist I

89.5000

2

0.70711

Ambulatory Surgery Coding Specialist

73.3333

3

6.4291

Total

84.7500

8

10.51190

7. Calculate the mean length of stay and the median age using Table 6A. Show your work.

Patient #

Age

Length of Stay (LOS)

Case 1

52.6

10

Case 2

81.9

12

Case 3

74.7

8

Case 4

23.8

4

Case 5

27.9

5

Case 6

42.5

6

Case 7

55.2

7

Case 8

77.0

9

Case 9

81.2

10

Case 10

53.8

8
Answer #1

7.

  • Mean is computed by dividing the sum of all the values by the total number of values. Arithmetic mean is represented by \bar{X} .

Formula :

\bar{X}= \frac{Sum\: of\, the\, values\left ( \sum x \right )}{Number \, of\, the\, values\left ( n \right )}

To calculate the mean length of stay,

\bar{X}= \frac{10 + 12 + 8 + 4 + 5 + 6 + 7 + 9 + 10 + 8}{10}

\bar{X}= \frac{79}{10}= 7.9

Hence, the mean of the length of stay is 7.9

  • Median of a set of values is the middle- most value when data is arranged in ascending order of magnitude.

Formula :  

M= \left ( \frac{n+1}{2} \right )^{th}\, observation\, or\, item

First, arrange the data in ascending order of magnitude.

23.8 , 27.9 , 42.5 , 52.6 , 53.8 , 55.2 , 74.7 , 77 , 81.2 , 81.9

M= \left ( \frac{10+1}{2} \right )^{th}\, observation\, or\, item

M= \left ( \frac{11}{2} \right )^{th}\, observation\, or\, item

M= 5.5 th observation or item

In ascending order the numbers are 23.8, 27.9, 42.5, 52.6, 53.8, 55.2, 74.7, 77, 81.2, 81.9 . There are 10 total numbers, so the 5th and 6th numbers are used to figure the median. (53.8 + 55.2) / 2 = 54.5

Hence, The median age is 54.5.